OpenStudy (anonymous):

what's the area between the curve y=x and y=xsinx from [0,pi/2]??

OpenStudy (anonymous):

mainly need help with setting up the equation...

OpenStudy (anonymous):

the most important part of this is figuring out which is larger on that interval

OpenStudy (anonymous):

since \(x\geq \sin(x)\) you get \[\int_0^{\frac{\pi}{2}}(x-\sin(x))dx\]